OpenStudy (anonymous):

Could someone help me figure out how to solve the rest of this? I need to find fog.. This is the initial equation: f(x)=x+7 and g(x)=3x+2 This is as far as I've gotten: fog = f(g(x)) = x(3x - 2)+7 Have I done this right so far? What do I do next? I have notes on this but I don't really understand how to do the next steps.

OpenStudy (anonymous):

I think I may have just figured it out....

zepdrix (zepdrix):

f composed of g.. hmm i think you're very close :) f(g(x)) So if g(x)=3x+2 then f(g(x))=f(3x+2) If f( ) = ( ) + 7 fill in the blank! :)

OpenStudy (anonymous):

Here's what I did: 1. f(x)=x+7 and g(x)=3x+2 2. fog=f(g(x))=x(3x-2)+7 3. fog=3x+2+7 (I switched from negative to positive/plus to minus) 4. Subtract 2 from itself and from 7 5. fog = 3x+5 Correct?

OpenStudy (noelgreco):

f(3x+2) means that you replace the x in function f with 3x+2

OpenStudy (noelgreco):

Also, in step 2 you've got 3x MINUS 2, not 3x+2

zepdrix (zepdrix):

f(x)= (x) + 7 f(x+y)= (x+y) + 7 f(potato)= (potato) + 7 whatever is inside of f, that's what you plug in for your original x. f(3x+2)= (3x+2) + 7 Does that make sense? :)